- My blog
- Post new blog entry
- Top bloggers
- Recent posts
Sort by:[Date]
UV_LED_Progress_Nov4_Pictures
Updated on Fri, 2016-11-04 09:25. Originally created by heppel on 2016-11-04 09:25.
Update on FMS LED/Frame progress
Updated on Thu, 2016-11-03 23:49. Originally created by bsumma on 2016-11-03 23:49.A brief update on the progress being made on the curing device for the FMS
See for reference to previous LED properties
Run 16 62 GeV d+Au Hit Errors
Updated on Tue, 2017-04-11 15:23. Originally created by witt on 2016-11-03 08:42.UPDATE: April 11, 2017
DAQ file used: st_physics_17141042_raw_5500001.daq
Run 16 20 GeV d+Au Hit Errors
Updated on Tue, 2017-04-11 15:18. Originally created by witt on 2016-11-03 08:38.UPDATE April 11, 2017
Run 16 200 GeV d+Au Hit Errors
Updated on Thu, 2016-11-03 08:31. Originally created by witt on 2016-11-03 08:31.DAQ file used: st_physics_17132063_raw_5000002.daq
Proceedings on Central Exclusive Production (DIffraction 2016)
Updated on Wed, 2016-11-16 17:08. Originally created by rafal_s on 2016-11-02 12:20.
Kielce 2016 v3
Updated on Wed, 2016-11-02 11:18. Originally created by siejka on 2016-11-02 11:18.Kielce 2016 v3
sPhenix_BtoD_simu
Updated on Sun, 2016-12-11 02:56. Originally created by xlchen on 2016-11-02 03:32.